342g of 20% by mass of Ba(oh)2 solution (specific gravity-0.57)is reacted with 1200ml of 2M HNO3 .IF the final density solution is same as pure water then molarity of the ion on resulting solution which decides the nature of the above solution is1).252).53)0.8884)none of these

Grade:12th pass

1 Answers

Paidimarrys Santhoshkumar
5 years ago
ANSWER IS C) 0.88M.
 
Ba(OH)2 Mass is 342 gr.
171gr is 1Mole
342gr is x
x=2 Moles
100% ------ 2moles
20% ............x
x=20*2/100=0.4moles
Ba(OH)Total moles 2*0.4=0.8
HNO3 Molarity is 2M
Total moles is M=n*1000/v
n=2*1200/1000=2.4moles
Total excess moles is acidic=2.4-0.8=1.6
Total volume of Ba(OH)2 is
d=w/v
v=342/0.57=600ml
HNO3 volume is 1200
total volume is 600+1200=1800
M=n*1000/v
M=1.6*1000/1800
M=0.88M